Hypothesis Testing I-sample means

1) Previous studies have shown that urban bus drivers have an extremely stressful job. Intervantion programs pput into place to imporve the conditions of bus drivers and their riders have been scheduled for implementation Prior to the intervention the following blood pressures were taken. Normal diastolic blood pressure averages 80 mm Hg.

Data: { 95, 58, 99, 81, 85, 81, 84, 77, 79, 100, 83, 89, 73, 79, 80, 90, 74, 79, 76, 65, 91,

88, 83, 90, 93, 69, 89, 95, 80, 70, 66, 94, 90,

77, 81, 75, 95, 73, 63 } N=40

Facts

Before you begin: be sure to descide on the distribution you'll use to test

Also, find appropriate mean and Sd

  1. Find the 90% confidence interval for mean diastolic blood pressure of the drivers

c. At the .10 = alpha (significance) level, determine if the data provide sufficient

evidence that the mean DBP of bus drivers exceeds the normal (80 mm hg).

Hypotheses:

Methodology/Calculations

Check conditions for the distribution you have chosen

Find critical value and test statistic

Conclusions: Complete Sentences
Answer #1

There are only 39 values given. So I do computation with these 39 values not n=40 Mean = x= Exi - 3189 = 81.77 39 n 50= {cx;-82:77 - 80 10 26 1439 1-077 Degrees of freedom = 0-1 = 38 critical value = t... 38 = 1. 304 Since calculated value is less th